Problem No. 2.4 0PS 2.xi-7x2-5x3 1 -7x1-32-9x3-3 Solve the system of linear equations by modifying it to REF and to RREF using elementary equivalent operations. Show REF and RREF of the system Show all your work, do not skip steps Displaying only answer is not enough to get credit Matrices may not be used
Answer #1

The augmented matrix for the given set of equations is

egin{bmatrix} 2 &-7 & -5 & 1 -4 &3 &-5 &-4 -7 & -3 & -9 &-3 end{bmatrix}

(a)

Row Echelon Form (REF)

R1 → R1/2

1 -7/2-5/2 1/2 -4 3-5-4

R_2 ightarrow 4R_1+R_2

1-7/2 -5/2 1/:2 0-11-15 -2

R_3 ightarrow7R_1+R_3

2 15 -2 0 11 0-55/2-53/2 2 1/2

R_2 ightarrow-R_2/11

0 -55/2 -53/2 1/2

R_3 ightarrow55R_2/2+R_3

egin{bmatrix} 1 &-7/2 & -5/2 & 1/2 0 &1 &15/11 &2/11 0 & 0 & 11 &11/2 end{bmatrix}

R_3 ightarrow R_3/11

egin{bmatrix} 1 &-7/2 & -5/2 & 1/2 0 &1 &15/11 &2/11 0 & 0 & 1 &1/2 end{bmatrix}

The transformed equation from the above REF matrix are

x_3=1/2

x_2+15x_3/11=2/11

Rightarrow x_2=2/11-15/11(1/2)=4/22-15/22=-11/22=-1/2

r1-7x212-513/2=1/2

Rightarrow x_1=1/2+7/2(-1/2)+5/2(1/2)=1/2-7/4+5/4=0

So the solution is x_1=0; x_2=-1/2 & x_3=1/2

(b)

Reduced Row Echelon Form (RREF)

R1 → R1/2

1 -7/2-5/2 1/2 -4 3-5-4

R_2 ightarrow 4R_1+R_2

1-7/2 -5/2 1/:2 0-11-15 -2

R_3 ightarrow7R_1+R_3

2 15 -2 0 11 0-55/2-53/2 2 1/2

R_2 ightarrow-R_2/11

0 -55/2 -53/2 1/2

R_3 ightarrow55R_2/2+R_3

egin{bmatrix} 1 &-7/2 & -5/2 & 1/2 0 &1 &15/11 &2/11 0 & 0 & 11 &11/2 end{bmatrix}

R_3 ightarrow R_3/11

egin{bmatrix} 1 &-7/2 & -5/2 & 1/2 0 &1 &15/11 &2/11 0 & 0 & 1 &1/2 end{bmatrix}

R_2 ightarrow -15R_3/11+R_2

egin{bmatrix} 1 &-7/2 & -5/2 & 1/2 0 &1 &0 &-1/2 0 & 0 & 1 &1/2 end{bmatrix}

R_1 ightarrow5R_3/2+R_1

egin{bmatrix} 1 &-7/2 & 0 & 7/4 0 &1 &0 &-1/2 0 & 0 & 1 &1/2 end{bmatrix}

R_1 ightarrow7R_2/2+R_1

2 01/ 001 010 100

From the above matrix, the solution to the given set of equations is x_1=0; x_2=-1/2 & x_3=1/2